what is the answer to s/6- 2/9=0
Found 2 solutions by Fombitz, Alan3354:
Answer by Fombitz(32388)
Answer by Alan3354(69443)

what is the answer to s/6- 2/9=0
------------
What is the question?
If you mean "What is the value of s?", then
s/6- 2/9=0
s/6 = 2/9
Multiply by 6
s = 12/9 = 4/3